OpenStudy (anonymous):

-0.6 + (-4a) = -1.4 bla. help.

OpenStudy (anonymous):

-.8 divided by -4 is .2

OpenStudy (anonymous):

So a = .2

OpenStudy (merchandize):

-0.6+(-4a)=-1.4 -4a=-1.4+0.6 -4a=-0.8 a=0.8/4 a=0.2

OpenStudy (anonymous):

It sometimes clarifies these problems to multiply by ten -6+-40a=-14